OpenStudy (anonymous):

CORRECT 4 MEDAL ! Lauren will make annual contributions in the amount of $4,770, on average, to a 401(k) over the next 34 years. Her employer will match 60% of her contributions. She is currently being taxed at 33%, but anticipates being taxed at 15% upon retirement. If her account grows at an average rate of 5.5% annually, what is the value of Lauren’s 401(k) upon retirement? $381,435.69 $395,228.47 <---- My answer $481,057.72 $610,297.11

OpenStudy (anonymous):

Total contributions (hers + employers) per year: 4770(1.60) = $7,632. FVoa = PMT [((1 + i)^n - 1) / i] FVoa = 7632[(1.055^34) - 1) / 0.055] = 7632[5.17424 / 0.055] = 7632(94.07712) = $717,996.5956, or $717,006.60 rounded Taxes don't come into play until she starts withdrawing the funds, so you can ignore them.

OpenStudy (anonymous):

thats wrong . Doesnt match nice try though from yahoo -.-

OpenStudy (anonymous):

Lol yp jst trying to help i really sck at math >.<

OpenStudy (solomonzelman):

I hate long and big word problems, b/c I can't really r=explain them, they are too long, but lets work it together, OK?

OpenStudy (anonymous):

ok :3

OpenStudy (solomonzelman):

Lauren will make annual contributions in the amount of $4,770, on average, to a 401(k) over the next 34 years. try to write this out.

OpenStudy (anonymous):

Do I use a formula for that o.o